Please help me with this. I am totally confused. Thanks.
1. Plot the graph of the equations 3x - 8y = 5 and 4x - 2y = 11 and interpret the result.
2. Plot the graph of the equations 4x - 6y = 2 and 2x - 3y = 1 and interpret the result.
3. Plot the graph of the equations 10x - 4y = 3 and 5x - 2y = 6 and interpret the result.
: Please help me with this. I am totally confused. Thanks.
1. Plot the graph of the equations 3x - 8y = 5 and 4x - 2y = 11 and interpret the result.
2. Plot the graph of the equations 4x - 6y = 2 and 2x - 3y = 1 and interpret the result.
3. Plot the graph of the equations 10x - 4y = 3 and 5x - 2y = 6 and interpret the result.

Answer by stanbon(19724)

You need two points to graph a line.
-------------------------
1. Plot the graph of the equations 3x - 8y = 5 and 4x - 2y = 11 and interpret the result.
3x-8y = 5
Let x = 0, then y = -4/8 giving you point (0,-5/8)
Let y = 0, then x = 5/3 givning you point (5/2, 0)
graph(400,300,-10,10,-10,10,(3/8)x - (5/8))
-------------------------------------------------------------
4x-2y = 11
Let x = 0, then y = -11/2 giving you point (0,-11/2)
Let y = 0, then x = 11/4 giving you point (11/4 , 0)
graph(400,300,-10,10,-10,10,2x -(11/2))
---------------------------------------------------------------
Now find the intersection of the two lines:
graph(400,300,-10,10,-10,10,(3/8)x - (5/8),2x -(11/2))
---------------------------------------------------------------
Find the solution which is the intersection point of the two lines.
=====================================================
Cheers,
Stan H.
Comment: Use the same procedure as above on your #2 and #3 problems.
2. Plot the graph of the equations 4x - 6y = 2 and 2x - 3y = 1 and interpret the result.
3. Plot the graph of the equations 10x - 4y = 3 and 5x - 2y = 6 and interpret the result.